# Sparse Portfolio Selection via Topological Data Analysis based Clustering

### v1

@@ 5.2 Strategy 1: Sparse Index Tracking
    IT with cardinality constraint: This portfolio corresponds to the optimization problem (15) with an additional constraint imposed on the maximum number of assets that can be selected within a tracking portfolio. Specifically, we impose the following restriction that the number of holdings with non-zero weights $\leq K_{max}$ where we set $K_{max}$ equal to the number of assets selected in the portfolio $C_{i},~{}i=1,2,\ldots,6$, to ensure a comparable number of asset for a fair comparison to the clustering based portfolios. Consequently, this yields six portfolios, denoted by $I_{i},~{}i=1,2,\ldots,6$. Given that the problem (15) with an additional constraint on the number of assets is a mixed-integer quadratic programming (MIQP) problem, it poses a challenging computational task. To tackle this computational bottleneck, we employ the Gurobi optimizer, which is recognized for its prowess in solving mixed-integer programming (MIP). For practical considerations, we follow Shu et al. (2020) and impose time limits of thirty minutes and one hour for solving each MIQP within every window. To distinguish, we denote the portfolios solved within a 30-minute time limit by $I_{i}^{(1)}$, while those solved with a one-hour time limit by $I_{i}^{(2)}$.
@@
Before analyzing the performance of the portfolios described above, we first examine the number of assets in the cluster containing the index resulting from APC based on the seven similarity matrices, $K_{i},~{}i=1,2,\ldots,7$, following our strategy 1. The mean and median number of assets in all windows are reported in Table 1. There are several observations. The number of clusters formed with TDA-based similarity matrices $K_{1}$ to $K_{4}$, as well as the correlation-based similarity matrices $K_{5}$ and $K_{6}$, is consistent over the sample period. A small number, on average, of assets in these clusters results in sparse portfolios as desired. For instance, the similarity matrix $K_{2}$ selects 40 assets (approximately 8%) on average, while the correlation-based matrix $K_{5}$ (or $K_{6}$) selects a further smaller number of assets. In contrast, the similarity matrix $K_{7}$, on average, selects almost 57% of assets, which is not desirable, and consequently, we drop the portfolio $C_{7}$ from our analysis. [^17] In Appendix 6.2, we present the study of the properties of the selected assets with respect to the benchmark index.
@@
Table 2 presents the out-of-sample performance of all portfolios in terms of five performance measures, i.e., correlation with the index (COR), tracking error (TE), excess mean return (EMR), information ratio (INFO), and the turnover ratio (TR). The best and second-best values among models are highlighted in bold and italics, respectively. Given that it is challenging to compare all the models simultaneously, we first identify the best-performing portfolio among those constructed based on clustering using TDA-based distance measures. Subsequently, we comprehensively compare these selected portfolios to those based on maximum similarity criteria and cardinality constraints.
@@
On the other hand, for $C_{2}$, the initial filtering step imparts stability to the problem, resulting in less frequent portfolio re-balancing and, consequently, lower transaction costs. In conclusion, no consensus exists for the best among $C_{2}$, $C_{4}$, and $I_{1t_{2}}-I_{6t_{2}}$. There is a trade-off between the tracking error and the transaction costs. Nonetheless, clustering-based portfolios require less computational time when compared to cardinality-constrained portfolios. To find the best portfolio, we further analyze their performance. That is, given the superior performance of the portfolios $C_{2}$, $C_{4}$ and $I_{1t_{2}}-I_{6t_{2}}$, we now compute the mean ratio of the in-sample tracking error and out-of-sample tracking error  [^18] We calculate the ratio of in-sample tracking error and out-of-sample tracking error in each window and then take average to find the mean Canakgoz and Beasley (2009). We also compare it to the mean ratio of the portfolio $TE_{all}$, which corresponds to a complete replication index tracking portfolio, i.e., a tracking portfolio constructed using all the assets constituting the index without penalty terms. A comparison of the mean ratios allows us to investigate the severity of the estimation error. Table 3 shows the average value of in-sample and out-of-sample tracking error across the windows and the average value of the corresponding ratios. A ratio of less than one across all portfolios indicates that, on average, there is a performance decline when transitioning from the in-sample period to the out-of-sample period. It is not surprising since the portfolio is optimized over the in-sample period, thus generating a smaller tracking error, while the same portfolio composition, when tracked over the out-of-sample period, the data that is unseen before. Nevertheless, it is remarkable that performance degradation is not excessively high, particularly for portfolios based on clustering techniques. This reaffirms the robustness of clustering-based portfolios in both the in-sample and out-of-sample periods.
@@
Remark 5.1.
To further validate the exceptional performance of our top-performing TDA-based sparse portfolio, $C_{2}$, we compare it to the tracking portfolios constructed using the co-integration technique, a methodology known for its effectiveness in index-tracking problems Alexander and Dimitriu (2005), Sant’Anna et al. (2017). We denote this portfolio by $CO$. The idea underlying the co-integration index tracking portfolio revolves around determining the optimal weights for a given set of assets $S_{1},\ldots,S_{n}$ that achieve the highest level of co-integration with the benchmark. Nevertheless, selecting the parameter $n$ is vital and demands careful consideration. For instance, to obtain a stable co-integrating vector of at least $n$ stocks out of $N$ stocks in the index, a mechanical search would entail examining an extensive array of portfolios - specifically, $N!/(N-n)!n!$ portfolios - as potential candidates for backtesting. However, in our paper, we adopt the approach presented in Sant’Anna et al. (2017) for constructing a co-integration-based index-tracking portfolio. More specifically, we start by randomly selecting 40 stocks, a choice motivated by our intention to compare it to the $C_{2}$ portfolio, which, on average, consists of 40 stocks. We run a regression analysis (for a comprehensive understanding of the co-integration model, please refer to Sant’Anna et al. (2017)) on the in-sample data and select this portfolio as a candidate if the residuals from the regression are stationary, as indicated by the Augmented Dickey-Fuller (ADF) test. The stationary residuals would ascertain the presence of co-integration between the index and asset price series. Otherwise, we discard the portfolio under consideration. We repeat this process for 100,000 regressions conducted sequentially for each portfolio. Each regression employs a unique combination of 40 assets. The final portfolio chosen is determined by the regression that yields the smallest sum of squared residuals. Our findings suggest that the $C_{2}$ portfolio outperforms the $CO$ portfolio in terms of all the performance metrics as reported in Table 4. For instance, the $C_{2}$ portfolio achieves a significantly lower tracking error of 5.14E-06 in contrast to 1.57E-05 achieved by the $CO$ portfolio. Similarly, the correlation of 0.968 with the index for the $C_{2}$ portfolio is significantly higher than 0.924 for the $CO$ portfolio. These findings suggest that TDA-based sparse index tracking portfolios perform better than traditional portfolios based on correlation and co-integration with the index.
@@
Covid-19 robustness check: As a robustness check, Table 5 shows the out-of-sample performance of tracking portfolios for data from the recent COVID-19 pandemic, July 2019 to mid-November 2020. Again, we use the daily prices of the S&P500 index and its 502 constituents as available in this period. We adopt the rolling window strategy with 21 days (approx one month) in-sample period and five days (approx one week) out-of-sample period. Among TDA-based portfolios, we observe that $C_{2}$ and $C_{4}$ still have the lowest and statistically significant tracking errors. Also, $C_{2}$ attains minimum turnover compared to other portfolios.
In contrast to the observations in the stable period, the maximum similarity-based portfolio $MS_{2}$ outperforms the best TDA-based portfolios, i.e., $C_{2}$ and $C_{4}$. For instance, $MS_{2}$ achieves a tracking error of 3.05E-05, better than 3.86E-05 and 3.53E-05 obtained by $C_{2}$ and $C_{4}$. Similarly, the turnover ratio 1.34 by $MS_{2}$ outperforms the ratio obtained by $C_{2}$ and $C_{4}$ of 1.66 and 1.94, respectively. The comparison of TDA-based portfolios against cardinality-based portfolios follows a pattern similar to the findings in the stable period in that there is no clear relation between them. However, considering the trade-off between the performance and the computational time, clustering-based portfolios outperform.
@@
TE 5.80E-05 2.91E-05 3.99E-05 3.27E-05 3.58E-05 3.34E-05
EMR 3.19E-04 -1.71E-04 4.09E-04 -2.49E-04 5.54E-04 4.73E-04
COR 9.29E-01 9.62E-01 9.51E-01 9.58E-01 9.55E-01 9.55E-01
INFO 4.18E-02 -3.17E-02 6.48E-02 -4.35E-02 9.27E-02 8.18E-02
TR 1.829 1.918 1.827 1.830 1.920 1.917
Case Study (Publicly available data for Index tracking model): For completeness, we also test the performance of our TDA-based sparse index tracking portfolios on the benchmark data sets from Beasley et al. (2003); Canakgoz and Beasley (2009) for index tracking problems. These data sets are drawn from several major world equity indices, and thus, analyzing the performance of these data sets gives additional evidence on the adequacy and effectiveness of the TDA-based portfolios.
@@ 8. Russell 3000 USA 3000 2151
We follow Beasley et al. (2003) and consider the first 145 weekly prices as in-sample observations and the following 146 weeks as out-of-sample periods to evaluate the performance of the models. We do not require a rolling window approach here, as there is only one training and testing period. Consequently, instead of reporting transaction costs, we find the Hinderfindahl index, a diversity index calculated as $\text{HHI}=\sum_{i=1}^{n}w_{i}^{2}.$ In the case of a naïve portfolio (which offers maximum diversification), the minimum value of HHI is obtained and is equal to $\frac{1}{n}$. Smaller values of HHI are desirable.
@@
Remark 5.2.
We also compare the performance of our best-performing TDA-based sparse portfolio, $C_{2}$, to the index tracking portfolio proposed in Beasley et al. (2003), which we denote by $C_{beas}$. It is worth noting that our TDA-based sparse portfolio, unlike $C_{beas}$, doesn’t require the number of assets to be included in the portfolio. Since the composition of the optimal index tracking portfolios in Beasley et al. (2003) is not publicly available, we directly use the tracking error values reported in that paper. Further, the tracking error used in Beasley et al. (2003) is
@@
Therefore, for a fair comparison, we also compute the tracking error, $TE_{beas}$, for our portfolio, $C_{2}$. Table 8 presents the out-of-sample $TE_{beas}$ for $C_{2}$ and $C_{beas}$ for data sets T1-T5, the indices used in Beasley et al. (2003). Interestingly, our $C_{2}$ portfolio outperforms $C_{beas}$ for all except T4 indices in terms of the tracking error $T_{beas}$.

### v2

@@ 5.2 Strategy 1: Sparse Index Tracking
    IT with cardinality constraint: This portfolio corresponds to the optimization problem (15) with an additional constraint imposed on the maximum number of assets that can be selected within a tracking portfolio. Specifically, we impose the following restriction that the number of holdings with non-zero weights $\leq K_{max}$ where we set $K_{max}$ equal to the number of assets selected in the portfolio $C_{i},~{}i=1,2,\ldots,6$, to ensure a comparable number of asset for a fair comparison to the clustering based portfolios. Consequently, this yields six portfolios, denoted by $I_{i},~{}i=1,2,\ldots,6$. Given that the problem (15) with an additional constraint on the number of assets is a mixed-integer quadratic programming (MIQP) problem, it poses a challenging computational task. To tackle this computational bottleneck, we employ the Gurobi optimizer, which is recognized for its prowess in solving mixed-integer programming (MIP). For practical considerations, we follow [77] and impose time limits of thirty minutes and one hour for solving each MIQP within every window. To distinguish, we denote the portfolios solved within a 30-minute time limit by $I_{i}^{(1)}$, while those solved with a one-hour time limit by $I_{i}^{(2)}$.
@@
Before analyzing the performance of the portfolios described above, we first examine the number of assets in the cluster containing the index resulting from APC based on the seven similarity matrices, $K_{i},~{}i=1,2,\ldots,7$, following our strategy 1. The mean and median number of assets in all windows are reported in Table 1. There are several observations. The number of clusters formed with TDA-based similarity matrices $K_{1}$ to $K_{4}$, as well as the correlation-based similarity matrices $K_{5}$ and $K_{6}$, is consistent over the sample period. A small number, on average, of assets in these clusters results in sparse portfolios as desired. For instance, the similarity matrix $K_{2}$ selects 40 assets (approximately 8%) on average, while the correlation-based matrix $K_{5}$ (or $K_{6}$) selects a further smaller number of assets. In contrast, the similarity matrix $K_{7}$, on average, selects almost 57% of assets, which is not desirable, and consequently, we drop the portfolio $C_{7}$ from our analysis. [^19] In Appendix C, we present the study of the properties of the selected assets with respect to the benchmark index.
@@
We also analyze the stability of cluster composition in time (across different windows) as well as among different similarity matrices, $K1$ to $K6$. To this end, we analyze the clustering structure over the period September 2020 to August 2022. We adopt a rolling window approach with a 6-month window size, shifting by one month, resulting in a total of 18 windows. We chose this recent period for its stable conditions and practical number of windows it offers, thus, making it easier to observe the variations in cluster composition and providing a reliable foundation for analysis. We start by calculating the Adjusted Rand Index (ARI) between the clustering composition in two consecutive windows for each similarity matrix. [^20] Refer Appendix for more details on ARI. We report the average value of ARI across the windows for each similarity matrix in Table 2. It is evident from the Table that each similarity matrix except $K4$ generates an average ARI above 0.9638 suggesting that these similarity matrices generate relative stable clusters over consecutive windows. In contrast, $K4$, with an average ARI of 0.8301, shows slightly lower stability, indicating that its clustering assignments may vary more significantly across different windows.
Table 2: Average Adjusted Rand Index (ARI) for affinity propagation clustering using $K_{1}-K_{6}$ across consecutive windows
Similarity Matrix $K1$ $K2$ $K3$ $K4$ $K5$ $K6$
Average ARI 0.9875 0.9638 0.9779 0.8301 0.9858 0.9857
Figure 5: A comparison of Jaccard similarity between the clusters, from different windows, containing the index for six kernel matrices.
Next, we compare the composition of the cluster containing the benchmark index, that is, the cluster selected for the index tracking problem. To do this, we compute the Jaccard similarity between these clusters from different windows, with results displayed in Figure 5.212121Refer Appendix for more details on Jaccard Similarity. Notably, we observe considerable variability in the similarity values across windows for each kernel, demonstrating how the cluster composition vary over time. There are several observations. First, the TDA based kernels, $K2$ to $K4$, show less similarity when compared to $K_{1}$ that produces relatively stable clusters. That is, the kernels $K_{2}$ to $K_{4}$ are able to capture the nuanced, non-linear relationships among assets, resulting in dynamic clusters that more accurately reflect market fluctuations in contrast to $K_{4}$ that is less responsive to market shifts. Second, the similarity matrix $K_{2}$ appears to strike a balance between stability and adaptability in the sense that the clusters in consecutive windows are neither the same nor completely changed, as evident by the first sub anti-diagonal (sub-diagonal above the main diagonal). [^22] Investors seeking a combination of responsiveness and stability may find $K_{2}$ particularly advantageous. The stability of $K_{2}$ is further supported by its lower turnover values compared to other TDA-based portfolios, underscoring its suitability for maintaining a stable investment strategy. Refer results section for more details. On the other hand, the kernels $K_{3}$ and $K_{4}$ shows significant change in the cluster composition, with exception by $K_{3}$ in the first half of the period. Third, the traditional similarity matrices, i.e., $K_{5}$ and $K_{6}$ also generate stable clusters, with $K_{5}$ demonstrating stability over a shorter period, while $K_{6}$ maintains stability in clusters of months (observe the small squares of different colors around the diagonal). Finally, observe the two prominent squares along the diagonal in the heat map of kernel $K_{2}$, the first one spanning the windows 1 to 12 and the second one covering the remaining six windows. These patterns corresponds to the bull (growth) and bear (fall) phase of the benchmark index (note that these clusters contain the benchmark) as illustrated in Figure 6. This separation is less distinct in the other similarity matrices, making $K_{2}$ particularly effective at highlighting shifts in asset behavior in response to market shifts. To conclude, the ability of the similarity matrix $K_{2}$ to capture the regime shifts advocates for the strength of the TDA-based approach in identifying nuanced market patterns.
Figure 6: Evolution of S&P500 over the period, September 2020 to August 202.
Finally, we compare the composition of the cluster containing the benchmark index across the six similarity matrices, $K_{1}$ to $K_{6}$. To this end, we compute the Jaccard similarity between clusters from each pair of kernels for each window, then take the average similarity across all windows. The results are presented in Figure 7. The relatively low similarity scores, ranging from 0 to 0.5, suggest that each similarity matrix captures distinct aspects of the data, generating clusters that differ significantly from one another, leading to diverse clusters. Notably, $K_{5}$ and $K_{6}$ show a moderate similarity compared to other pairs since they both try to capture the correlation among the assets. When comparing TDA-based similarity matrices to correlation-based matrices, $K_{2}$ shows the highest similarity to both $K_{5}$ and $K_{6}$. This suggests that TDA-based approaches could provide a valuable alternative to traditional methods, capturing not only the correlation but potentially provide deeper and robust insights into the asset relationships, particularly when dealing with non-linear dependencies.
Figure 7: Average similarity (Jaccard) across the cluster composition (containing the benchmark index) across the six similarity matrices, $K_{1}$ to $K_{6}$.
@@
Table 3 presents the out-of-sample performance of all portfolios in terms of five performance measures, i.e., correlation with the index (COR), tracking error (TE), excess mean return (EMR), information ratio (INFO), and the turnover ratio (TR). The best and second-best values among models are highlighted in bold and italics, respectively. The average computation time (from data to portfolio weights) per rolling window for portfolios $C_{1}$ to $C_{4}$ are 10.36, 15.09, 11.46, and 16.15 minutes, respectively. In contrast, the computational time for portfolios $C_{5}$ and $C_{6}$, based on correlation coefficients, are 1.5 and 2.47 minutes. The time for $I_{t_{1}}$ and $I_{t_{2}}$ are respectively 30 minutes and 60 minutes by definition of these portfolios. The computational time for $C_{1}$ to $C_{4}$ is relatively larger since these portfolios involves calculating the Wasserstein distance between persistence diagrams. While TDA based portfolios take more time, these durations are manageable given the fact the obtained portfolios are rebalanced on monthly basis, thus, making the computational load relatively insignificant. Given that it is challenging to compare all the models simultaneously, we first identify the best-performing portfolio among those constructed based on clustering using TDA-based distance measures. Subsequently, we comprehensively compare these selected portfolios to those based on maximum similarity criteria and cardinality constraints.
@@
On the other hand, for $C_{2}$, the initial filtering step imparts stability to the problem, resulting in less frequent portfolio re-balancing and, consequently, lower transaction costs. In conclusion, no consensus exists for the best among $C_{2}$, $C_{4}$, and $I_{1t_{2}}-I_{6t_{2}}$. There is a trade-off between the tracking error and the transaction costs. Nonetheless, clustering-based portfolios require less computational time when compared to cardinality-constrained portfolios. To find the best portfolio, we further analyze their performance. That is, given the superior performance of the portfolios $C_{2}$, $C_{4}$ and $I_{1t_{2}}-I_{6t_{2}}$, we now compute the mean ratio of the in-sample tracking error and out-of-sample tracking error  [^23] We calculate the ratio of in-sample tracking error and out-of-sample tracking error in each window and then take average to find the mean [15]. We also compare it to the mean ratio of the portfolio $TE_{all}$, which corresponds to a complete replication index tracking portfolio, i.e., a tracking portfolio constructed using all the assets constituting the index without penalty terms. A comparison of the mean ratios allows us to investigate the severity of the estimation error. Table 4 shows the average value of in-sample and out-of-sample tracking error across the windows and the average value of the corresponding ratios. A ratio of less than one across all portfolios indicates that, on average, there is a performance decline when transitioning from the in-sample period to the out-of-sample period. It is not surprising since the portfolio is optimized over the in-sample period, thus generating a smaller tracking error, while the same portfolio composition, when tracked over the out-of-sample period, the data that is unseen before. Nevertheless, it is remarkable that performance degradation is not excessively high, particularly for portfolios based on clustering techniques. This reaffirms the robustness of clustering-based portfolios in both the in-sample and out-of-sample periods.
@@
Remark 3.
To further validate the exceptional performance of our top-performing TDA-based sparse portfolio, $C_{2}$, we compare it to the tracking portfolios constructed using the co-integration technique, a methodology known for its effectiveness in index-tracking problems [5], [73]. We denote this portfolio by $CO$. The idea underlying the co-integration index tracking portfolio revolves around determining the optimal weights for a given set of assets $S_{1},\ldots,S_{n}$ that achieve the highest level of co-integration with the benchmark. Nevertheless, selecting the parameter $n$ is vital and demands careful consideration. For instance, to obtain a stable co-integrating vector of at least $n$ stocks out of $N$ stocks in the index, a mechanical search would entail examining an extensive array of portfolios - specifically, $N!/(N-n)!n!$ portfolios - as potential candidates for backtesting. However, in our paper, we adopt the approach presented in [73] for constructing a co-integration-based index-tracking portfolio. More specifically, we start by randomly selecting 40 stocks, a choice motivated by our intention to compare it to the $C_{2}$ portfolio, which, on average, consists of 40 stocks. We run a regression analysis (for a comprehensive understanding of the co-integration model, please refer to [73]) on the in-sample data and select this portfolio as a candidate if the residuals from the regression are stationary, as indicated by the Augmented Dickey-Fuller (ADF) test. The stationary residuals would ascertain the presence of co-integration between the index and asset price series. Otherwise, we discard the portfolio under consideration. We repeat this process for 100,000 regressions conducted sequentially for each portfolio. Each regression employs a unique combination of 40 assets. The final portfolio chosen is determined by the regression that yields the smallest sum of squared residuals. Our findings suggest that the $C_{2}$ portfolio outperforms the $CO$ portfolio in terms of all the performance metrics as reported in Table 5. For instance, the $C_{2}$ portfolio achieves a significantly lower tracking error of 5.14E-06 in contrast to 1.57E-05 achieved by the $CO$ portfolio. Similarly, the correlation of 0.968 with the index for the $C_{2}$ portfolio is significantly higher than 0.924 for the $CO$ portfolio. These findings suggest that TDA-based sparse index tracking portfolios perform better than traditional portfolios based on correlation and co-integration with the index.
@@
Covid-19 robustness check: As a robustness check, Table 6 shows the out-of-sample performance of tracking portfolios for data from the recent COVID-19 pandemic, August 2019 to August 2020. Again, we use the daily prices of the S&P500 index and its 502 constituents as available in this period. We adopt the rolling window strategy with 21 days (approx one month) in-sample period and five days (approx one week) out-of-sample period. Among TDA-based portfolios, we observe that $C_{2}$ and $C_{4}$ still have the lowest and statistically significant tracking errors. Also, $C_{1}$ attains minimum turnover while $C_{2}$ attains second minimum turnover compared to other portfolios, an observation that aligns with observations in Figure 5 where we observed that $C_{1}$ generates the most similar clusters.
In contrast to the stable period, the best performing TDA-based portfolios, i.e., $C_{2}$ and $C_{4}$ outperforms the maximum similarity-based portfolio $MS_{2}$ outperforms in terms of tracking error and EMR, but falls short in terms of correlation and turnover. Specifically, $C_{2}$ and $C_{4}$ achieves tracking errors of 2.43E-03, and 2.56E-03, respectively, which is better than 4.16E-03 recorded by $MS_{2}$. However, $MS_{2}$ shows a favorable turnover ratio of 0.77, compared to 1.39 and 1.98 for $C_{2}$ and $C_{4}$ respectively. This is likely because $MS_{2}$ maintains a fixed portfolio size of 20 assets, thereby capping the maximum number of position changes at 20. The comparison of TDA-based portfolios against cardinality-based portfolios shows that those based on the similarity matrix $K_{2}$ outperforms other cardinality based portfolios. Additionally, TDA based portfolio $C_{2}$ outperforms the best cardinality based portfolios, $I_{2t_{1}}$ and $I_{2t_{2}}$ by achieving a lower tracking error and turnover, alongside higher correlation. Therefore, considering the trade-off between the performance and the computational time, clustering-based portfolios emerge as the superiod alternative.
@@
TE 6.84E-03 3.22E-03 6.05E-03 3.24E-03 6.79E-03 5.64E-03
EMR -5.98E-04 1.58E-04 -2.62E-04 -9.55E-05 -2.85E-04 3.98E-06
COR 9.72E-01 9.93E-01 9.77E-01 9.93E-01 9.68E-01 9.80E-01
INFO -8.75E-02 4.90E-02 -4.33E-02 -2.95E-02 -4.20E-02 7.05E-04
TR 1.63 1.50 1.41 1.50 1.68 1.77
Case Study (Publicly available data for Index tracking model): For completeness, we also test the performance of our TDA-based sparse index tracking portfolios on the benchmark data sets from [9, 15] for index tracking problems. These data sets are drawn from several major world equity indices, and thus, analyzing the performance of these data sets gives additional evidence on the adequacy and effectiveness of the TDA-based portfolios.
@@ 8. Russell 3000 USA 3000 2151
We follow [9] and consider the first 145 weekly prices as in-sample observations and the following 146 weeks as out-of-sample periods to evaluate the performance of the models. We do not require a rolling window approach here, as there is only one training and testing period. Consequently, instead of reporting transaction costs, we find the Hinderfindahl index, a diversity index calculated as $\text{HHI}=\sum_{i=1}^{n}w_{i}^{2}.$ In the case of a naïve portfolio (which offers maximum diversification), the minimum value of HHI is obtained and is equal to $\frac{1}{n}$. Smaller values of HHI are desirable.
@@
Remark 4.
We also compare the performance of our best-performing TDA-based sparse portfolio, $C_{2}$, to the index tracking portfolio proposed in [9], which we denote by $C_{beas}$. It is worth noting that our TDA-based sparse portfolio, unlike $C_{beas}$, doesn’t require the number of assets to be included in the portfolio. Since the composition of the optimal index tracking portfolios in [9] is not publicly available, we directly use the tracking error values reported in that paper. Further, the tracking error used in [9] is
@@
Therefore, for a fair comparison, we also compute the tracking error, $TE_{beas}$, for our portfolio, $C_{2}$. Table 9 presents the out-of-sample $TE_{beas}$ for $C_{2}$ and $C_{beas}$ for data sets T1-T5, the indices used in [9]. Interestingly, our $C_{2}$ portfolio outperforms $C_{beas}$ for all except T4 indices in terms of the tracking error $T_{beas}$.